OpenStudy (anonymous):

9c-4=3c-8 i get that its 6c-4=8 but is it an 8 or a -8

OpenStudy (anonymous):

add 4 to each side 6c=12 divide each side by 6 c is 2

OpenStudy (anonymous):

okay but this is a review prob for my test tomorrow and i was wondering if it would be -8 or 8

OpenStudy (anonymous):

as in answer? well plus in 8 or -8 in the original equation. it'll work if they equal each other. try it

OpenStudy (anonymous):

6c +4=8 or -8

OpenStudy (anonymous):

9c-4=3c-8 isolate for c 6c-4=-8 add 4 to each side 6c=-4 divide both sides by 6 c=-4/6